OpenStudy (anonymous):

Select the point that is on the right side of the vertex of y = |x + 3| + 4.

OpenStudy (anonymous):

I don't understand, you want vetex?

OpenStudy (anonymous):

y = |x + 3| + 4 y = ± (x + 3) + 4 If y = (x + 3) + 4, y = x + 7 If y = - (x + 3) + 4, y = - x - 3 + 4 y = - x + 1 y = y, so x + 7 = - x + 1 x + x = 1 - 7 2x = - 6 x = - 6 / 2 x = - 3

OpenStudy (anonymous):

-3,0..........->>>>>>>>>> then -3,4 <<<<<<<this is the answer

OpenStudy (anonymous):

and y = - 3 + 7 y = 4

OpenStudy (anonymous):

thank you